OpenStudy (anonymous):

Can you add the square root of 6 and the square root of 24?

OpenStudy (phi):

by putting a + between them if you simplify things, you can change the form

OpenStudy (anonymous):

\[\sqrt{6}+\sqrt{24}=\sqrt{6}+\sqrt{4*6}=\sqrt{6}+2\sqrt{6}=3\sqrt{6}\]

OpenStudy (raden):

use the property : m sqrt(a) + n sqrt(a) = (m+n)sqrt(a)